What is the average deviation of the following numbers:

2.00, 2.48, 1.90
Answer #1

Mean of the values ( ) = (2.00+2.48+1.90)/3 = 2.127.

Average deviation (s) =

n is number of readings.

Hence

S = |2.127 -2| +|2.127-2.48| + |2.127 -1.90|/3

S = (0.127 + 0.353 + 0.227)/3

S = 0.235

So average deviation is 0.235.